OpenStudy (anonymous):

What is the mean of the following set of data? Round your answer to two decimal places. 34, 56, 22, 49, 37, 20, 61, 48, 22

Mehek (mehek14):

1) add all the numbers : 35 + 56 + 22 + 49 + 37+ 20 + 61 + 48 + 22 = 350 2) Divide 350 by 9 3) 38.89 (rounded answer)

OpenStudy (anonymous):

\[=\frac{ 34+56+22+49+37+20+61+48+22 }{ 9 }\]\[=\frac{ 349 }{ 9 }\]\[=?\] SOLVE IT @BradSuga

OpenStudy (anonymous):

it's not 39.9 @heavenfun13

OpenStudy (anonymous):

38.78

OpenStudy (anonymous):

38.8

OpenStudy (anonymous):

yup it's 38.78 :)

OpenStudy (anonymous):

Would this be the same way? Find the median of the following set of data. 13, 11, 4, 5, 6, 9, 10, 12, 15, 16Find the median of the following set of data. 13, 11, 4, 5, 6, 9, 10, 12, 15, 16

OpenStudy (anonymous):

Oops pasted it twice

OpenStudy (anonymous):

it will not be the same way

OpenStudy (anonymous):

Any Idea how @MARC_

OpenStudy (anonymous):

First arrange it in a numerical order @BradSuga

OpenStudy (anonymous):

From the smallest number to the biggest number

OpenStudy (anonymous):

\[4,5,6,9,10,11,12,13,15,16\]

OpenStudy (anonymous):

 @BradSuga